### v2

# SPELL: Self-Play Reinforcement Learning for evolving Long-Context Language Models

Abstract
Progress in long-context reasoning for large language models (LLMs) has lagged behind other recent advances. This gap arises not only from the intrinsic difficulty of processing long texts, but also from the scarcity of reliable human annotations and programmatically verifiable reward signals. In this paper, we propose SPELL, a multi-role self-play reinforcement learning framework that enables scalable, label-free optimization for long-context reasoning. SPELL integrates three cyclical roles—questioner, responder, and verifier—within a single model to enable continual self-improvement. The questioner generates questions from raw documents paired with reference answers; the responder learns to solve these questions based on the documents; and the verifier evaluates semantic equivalence between the responder’s output and the questioner’s reference answer, producing reward signals to guide continual training. To stabilize training, we introduce an automated curriculum that gradually increases document length and a reward function that adapts question difficulty to the model’s evolving capabilities. Extensive experiments on six long-context benchmarks show that SPELL consistently improves performance across diverse LLMs and outperforms equally sized models fine-tuned on large-scale annotated data. Notably, SPELL achieves an average 7.6-point gain in pass@8 on the strong reasoning model Qwen3-30B-A3B-Thinking, raising its performance ceiling and showing promise for scaling to even more capable models.
1 Introduction
In recent years, reinforcement learning (RL) has emerged as a promising approach for enhancing the reasoning capabilities of large language models (LLMs) (guo2025deepseek; qwen3_technical_report; openai2024openaio1card; team2025kimi). Among these methods, reinforcement learning with verifiable rewards (RLVR) has shown particular promise in domains where correctness can be programmatically verified, such as mathematics, logical reasoning, and software engineering (lambert2024tulu; OpenReasonerZero2025; liu2025synlogic; wei2025swe). RLVR methods employ rule-based or programmatic verifiers to generate reward signals, which then guide policy optimization through algorithms such as Proximal Policy Optimization (PPO) (schulman2017proximal), Group Relative Policy Optimization (GRPO) (shao2024deepseekmath), and related variants (shao2024deepseekmath; yue2025vapo; liu2025understanding).
Despite these advances, most RLVR research has been restricted to short-context settings (e.g., <1024 tokens), where models primarily rely on their parametric knowledge for reasoning (qwenlongl1). In contrast, reasoning over long documents like long-context question answering requires not only locating relevant evidence scattered across extended contexts but also executing multi-step reasoning. Extending RLVR to long-context reasoning presents significant challenges, which stem from the inherent difficulty of processing long texts, as well as two critical bottlenecks: the prohibitive cost and unreliability of human annotations, and the absence of programmatically verifiable rewards.
Empirical evidence highlights the severity of these issues. On benchmarks such as LongBench-V2, human accuracy for extra-long multiple-choice reasoning tasks drops to 25.1% — effectively approaching random chance (longbench_v2). This not only limits the performance achievable under human supervision but also imposes a scalability ceiling, particularly as LLMs approach superhuman reasoning capabilities (arz). Specifically, as context length grows, producing reliable annotations becomes increasingly costly and unstable, and supervision diversity diminishes. Moreover, the lack of verifiable reward mechanisms in long-context settings further constrains the applicability of RLVR, posing a fundamental challenge to advancing reasoning capabilities at scale.
To address these limitations, we turn to self-play RL, where a single model learns to self-evolve by generating and solving its own tasks without human labels (zhou2025selfchallenginglanguagemodelagents; selfquestion; rzero). However, applying self-play to long-context reasoning poses a unique challenge: answers may be semantically correct yet differ substantially in expression, rendering string matching or naive majority voting unreliable reward signals. Thus, the model should not only generate questions and answers, but also verify its own solutions reliably. This observation motivates our framework, in which one LLM assumes three complementary roles: questioning, responding, and verifying.
In this paper, we introduce SPELL (Self-Play Reinforcement Learning for Evolving Long-Context Language Models), a self-play RL framework for long-context reasoning. In this setup, a unified policy alternates among three roles: the questioner, which formulates questions with reference answers from raw documents; the responder, which attempts to solve them; and the verifier, which compares the responder’s output with the reference answer to produce reward signals for joint optimization. To steer this process, SPELL incorporates three key design elements. First, a verifier trained for self-consistency on verifiable tasks produces stable rewards, even for outputs that cannot be verified by strict rules, thereby overcoming the brittleness of string matching. Second, an automated curriculum uses a history memory of question–answer pairs and documents to progressively increase task difficulty. A Gaussian-shaped reward further calibrates difficulty around the responder’s competence frontier, ensuring questions are neither too easy nor impossibly difficult. Third, a role-specific dynamic sampling strategy balances contributions across roles to stabilize training of the shared policy. Together, these components form a self-sufficient, closed-loop system that enables LLMs to autonomously evolve long-context reasoning without human-labeled data
We evaluate SPELL across 12 open-source LLMs ranging from 4B to 32B parameters, including both dense and Mixture-of-Experts (MoE) architectures. On six long-context QA benchmarks, SPELL delivers consistent performance gains. Remarkably, training a base model with SPELL enables it to surpass its instruction-tuned counterpart that relies on extensive human-annotated data, highlighting the data efficiency of our label-free self-play approach. Against a strong RLVR baseline trained on a static dataset synthesized by DeepSeek-R1-0528 (guo2025deepseek), SPELL achieves larger and more reliable gains. For capable models such as Qwen3-30B-A3B-Thinking, SPELL’s dynamic curriculum continually elevates performance and enables it to outperform the leading gemini-2.5-pro (comanici2025gemini) in pass@4. These findings firmly establish our self-play approach as a scalable and effective path toward advanced long-context reasoning without human supervision.
2 Related Work
2.1 Long-Context Alignment
Developing long-context language models (LCLMs) has become a central research area, as many real-world applications require reasoning over extended inputs (liu2025comprehensivesurveylongcontext). A dominant paradigm in this field is to enhance models through various post-training alignment techniques on well-curated, synthesized datasets. One prominent approach is supervised fine-tuning (SFT). For instance, LongAlign (bai2024longalign) utilizes a self-instruct pipeline to construct a large-scale, long-instruction dataset for SFT, while MIMG (chen2024essential) employs a multi-agent system to generate more complex, multi-hop reasoning data. Another line of work focuses on preference optimization. LongPO (longpo) and SoLoPO (solopo) generate preference pairs by contrasting outputs from compressed versus full contexts and leverage direct preference optimization (DPO) (rafailov2023direct) to transfer short-context capabilities to longer inputs. More recently, QwenLong-L1 (qwenlongl1) introduces the concept of long-context reasoning RL and employs a two-stage progressive context scaling training to develop long-context reasoning models. While proven effective, all these approaches exhibit some form of reliance on external supervision.
2.2 Self-Play Language Models
To mitigate the reliance on external supervision, such as human annotation or labeled datasets, researchers are developing self-play language models (gao2025survey). These models achieve autonomous improvement by generating their own training data, reward signals, or both. These approaches can be categorized into two paradigms: multi-model optimization, which co-evolves several distinct models, and single-model optimization, where one model assumes multiple roles.
Multi-Model Optimization
This paradigm orchestrates the co-evolution of multiple specialized models. In Mutual-Taught (mutualtaught), a policy model and a reward model are reciprocally and iteratively refined: the policy model generates data to enhance the reward model, which in turn provides more accurate feedback to improve the policy model. Similarly, in R-Zero (rzero), a challenger and a solver LLM are independently optimized and co-evolve: the challenger generates new, challenging math problems for which it is rewarded based on the solver’s uncertainty, and the solver is fine-tuned on these questions, rewarded for correctly solving them. While effective, this paradigm substantially increases systemic complexity, and performance gains often plateau after a finite number of iterations.
Single-Model Optimization
In contrast to the multi-model approach, single-model optimization reduces systemic complexity by leveraging a single model to assume multiple roles for self-improvement. For instance, Absolute Zero Reasoner (AZR) (arz) employs one model as both a proposer that generates complementary coding tasks (induction, abduction, and deduction) and a solver that addresses them, with code execution feedback serving as the reward signal. Similarly, the Self-Challenging Agent (SCA) (zhou2025selfchallenginglanguagemodelagents) employs a model to formulate novel “Code-as-Task” questions and subsequently solve them, using self-generated code functions to provide the verification signal. A significant limitation of such approaches, however, is their dependence on external code executors, which confines their applicability to domains with programmatically verifiable outcomes. To overcome this limitation and enhance autonomy, other approaches further incorporate self-rewarding mechanisms that leverage self-consistency (zuo2025ttrl), internal confidence scores (li2025confidence), or self-generated evaluations (yuan2025selfrewardinglanguagemodels). For example, Self-Questioning Language Models (SQLM) (selfquestion) utilize a model to propose and then answer questions, adapting its reward mechanism between self-consistency for arithmetic and proposer-generated unit tests for coding. Similarly, the Self-Rewarding Self-Improving framework (simonds2025selfrewardingselfimproving) also generates its own questions and solutions but uses a self-judging mechanism for reward computation.
Our work extends the single-policy self-play paradigm to long-context understanding and reasoning. Unlike existing methods that focus on short-context tasks like coding or math, SPELL is designed for reasoning over long documents. In our framework, a single LLM learns by playing three roles: a questioner, a responder, and a verifier. These roles interact to create a self-sufficient learning loop for long-context comprehension, thereby addressing a key gap in current self-play learning research.
3 Preliminaries
Long-Context Reinforcement Learning
We formulate the long-context generation task as a reinforcement learning (RL) problem. Given a set of documents and a question , the goal of long-context RL is to optimize a policy model to generate a response that maximizes a reward function . The standard objective is to maximize the KL-regularized expected reward (schulman2017equivalence; qwenlongl1):
where , is the training dataset, denotes a reference policy, and controls the strength of the KL regularization to prevent large deviations from the reference policy.
Group Relative Policy Optimization (GRPO)
For long-context inputs, the quadratic complexity of the attention mechanism renders PPO (schulman2017proximal), which relies on generalized advantage estimation (GAE) (schulman2015high) via a value network, computationally prohibitive. Therefore, we employ GRPO (shao2024deepseekmath) to optimize the objective in Eq. (1). For each input , GRPO first samples a group of candidate responses from the old policy . It then estimates the advantage through group-wise reward z-score normalization, thereby obviating the need for a separate value network. Formally, the objective is:
where is the importance sampling ratio for token in sequence . The group-relative advantage is shared across tokens of the -th sequence and computed by normalizing the sequence-level rewards :
4 The SPELL framework
In this section, we detail the core design of SPELL, a self-play reinforcement learning framework that enables LLMs to improve their long-context reasoning capabilities without external supervision. The key principle of SPELL is that a single policy model dynamically assumes three complementary roles: a questioner (), a responder (), and a verifier (). Through their interaction, the model autonomously generates and solves questions while producing reliable reward signals. This closed-loop interaction creates an evolving curriculum in which the model progressively adapts to longer contexts and more complex reasoning (Section 4.1). Role-specific reward designs (Section 4.2) and a unified optimization procedure (Section 4.3) jointly drive this co-evolution.
4.1 The Self-Play Evolutionary Loop
As illustrated in Figure 2 and Algorithm 1, SPELL proceeds iteratively: given a cluster of documents and a task type111Details of dataset construction and task definition are provided in Appendix C.1. , the policy first generates new questions,222To direct the policy in enacting three distinct roles, we adopt zero-shot prompting using tailored templates for each role and task type. Details of these templates are provided in Appendix E. then attempts to solve them, and finally verifies the solutions before performing a unified policy update.
Questioning
The questioner generates new question–answer pairs in an iterative curriculum. In the very first iteration, it is conditioned only on a randomly sampled subset of documents () and produces a pair . After each solvable pair is created, we append it to a history memory that stores the most recent solvable question–answer pairs and their associated source documents: . In subsequent iterations, the questioner is conditioned on both a newly sampled subset and the stored memory. The resulting context is . As the memory fills, the context for the questioner expands to include both previously seen and newly sampled documents, which allows the questioner to generate questions that integrate information across more documents. The history memory also raises difficulty by including past : these exemplars discourage redundancy and, via prompting, push to generate harder questions than those already solved. Consequently, the questioner’s difficulty increases for two complementary reasons: (1) the context expands over iterations as more documents are brought into scope, and (2) explicit conditioning on historical encourages the model to escalate question complexity.
Responding
The responder attempts to solve the generated question based on documents. To mitigate the generation of non-grounded or hallucinated questions, we employ a grounding filter process to discard questions that can be trivially answered without documents. For valid questions, the responder is presented with the complete set of documents, where the remaining documents unseen by the questioner serve as distractors to increase grounding and reasoning difficulty. This design enforces reliance on the provided document context rather than parametric memory. To encourage exploration of diverse reasoning trajectories, the responder generates independent rollouts .
Verifying
The verifier evaluates the semantic equivalence between the responder’s output and the questioner’s reference answer . For each , it produces independent binary judgments , which are then aggregated through majority voting:
where is the indicator function. This ensemble-based verification reduces variance and produces a stable, semantically aware reward signal, which is essential for sustaining a self-play system. . This step forces the questioner to create questions that are deeply tied to the provided text, thereby targeting the grounding ability of the responder.
4.2 Role-Specific Reward Design
The three roles co-evolve under specialized rewards that align their objectives while remaining compatible within a single shared policy. In what follows, we detail these rewards.
Verifier
The verifier is trained to improve its judgment reliability through self-consistency (wang2022self; zuo2025ttrl). For a candidate output , the verifier produces rollouts with judgments . Each rollout is then assigned a reward:
where is the majority vote over rollouts.
Responder
The responder’s reward for the -th solution is the maximum of a deterministic, rule-based check and the verifier’s consensus score, denoted as:
The rule-based function, , provides a binary reward based on cover exact match (CEM) criteria (qwenlongl1; r1searchplusplus)—it returns 1 if the ground-truth answer appears in the generated response and 0 otherwise. The maximum reward plays a crucial role: when is a correct paraphrase that CEM fails to capture, a majority vote of prevents the policy from being misled by false-negative noise, which stabilizes learning and encourages continual improvement.
Questioner
The questioner is incentivized to generate questions of intermediate difficulty, as learning is most efficient at the frontier of the LLM’s capabilities (bae2025online; rzero). For binary-reward tasks, this frontier corresponds to a success probability of 0.5, which maximizes reward variance and provides the richest learning signal. We therefore define the questioner’s reward as a Gaussian function centered at this optimal point. Given the responder’s average success rate, , the reward is:
We set the mean to target the point of maximum learning efficiency and the standard deviation to concentrate the reward around this level. Additionally, the questioner is penalized for producing ill-formatted (e.g., non-parsable) question–answer pairs or questions that can be solved without context, thereby enforcing both correct formatting and strong grounding in the provided text.
4.3 Unified Policy Optimization
A central feature of SPELL is that samples generated under different roles supervise a single policy . The optimization must control both sample efficiency and gradient balance across roles.
Role-Specific Dynamic Sampling
The raw samples collected for each document instance are highly imbalanced: one questioner sample, responder samples, and verifier judgments. To prevent the verifier’s samples from dominating updates and to prioritize improvements in the responder’s document-grounded reasoning, we introduce a role-specific sampling strategy that leverages the statistical structure of each role’s signals. For the responder, we retain all groups with non-zero reward variance (). The associated questions are labeled as positives for the questioner, and an equal number of negatives are drawn from questions with non-positive reward, as defined in Eq. (7). For the verifier, we preserve instances where the majority vote agrees with the rule-based check and subsample groups with conflicting verifications to match the number of questions. This role-specific sampling strategy reduces the training set to roughly of all samples, accelerates optimization, and prevents the responder’s gradients from being overwhelmed by verifier samples. Importantly, although most verifier samples are omitted, their collection cost is low, see Appendix D.1.
Advantage Estimation
For the responder and verifier, which generate outputs per prompt, we use group-level advantage estimation as defined in Eq. (3):
The questioner generates only a single output per instance and thus lacks a group-level baseline. Therefore, we adapt the normalization method from REINFORCE++-baseline (hu2025reinforce++) and normalize its reward against other questioner rewards within the training batch :
Unified Policy Update
After collecting and sampling a batch of samples, the policy parameters are updated by jointly optimizing the GRPO objective across all three roles:
The updated is reused to execute all roles in the next iteration. This closes the self-evolutionary cycle and keeps one unified policy for questioning, responding, and verifying.
5 Experiments
5.1 Experimental Setup
Training Details
Our SPELL RL framework is implemented using VeRL (verl). During generation, we employ a sampling temperature of 0.7 and a top- value of 0.95. The maximum input length is 16K tokens, while the maximum output length is set to 4K for non-reasoning models and extended to 20K tokens for reasoning models. To balance rollout diversity and computational efficiency, we utilize a group size of . The maximum number of recent solvable question–answer pairs cached in history memory is set to , and the number of candidate documents drawn when proposing a new question is set to . We conduct a purely on-policy RL training with a batch size of 128 and a constant learning rate of . At the beginning of each rollout, we randomly sample one of three predefined task formats—document general QA, financial math QA, or multiple-choice—along with a relevant document list from the corpus. Prompt templates for each task and each role are provided in Appendix E. For the RLVR baseline, we synthesize a dataset using DeepSeek-R1-0528 (guo2025deepseek) over the same document corpus and maintain identical hyperparameters to ensure a fair comparison. For comprehensive details on data construction, RL algorithm, and baselines, please refer to Appendix C.1, C.3, and C.4.
Evaluation Benchmarks
We evaluate our models on six long-context benchmarks, spanning multiple-choice QA on LongBench-V2 (longbench_v2) and multi-hop QA across Frames (krishna2024fact), HotpotQA (yang2018hotpotqa), 2WikiMultihopQA (ho2020constructing), MuSiQue (trivedi2022musique)333We use the subsets of HotpotQA, 2WikiMultihopQA, and MuSiQue from LongBench (longbench_v1)., and the DocMath (zhao2024docmath) for financial report reasoning task. We evaluate all models with maximum input lengths of 16K and 100K tokens, and report the average accuracy over eight runs. Further details on the benchmarks and evaluation protocol are available in Appendix C.2.
| Models | 16K | 100K | ||||||||
| DocMath | Frames | LB-MQA | LB-V2 | Avg. | DocMath | Frames | LB-MQA | LB-V2 | Avg. | |
| Base Models | ||||||||||
| Qwen2.5-7B | 10.9 | 27.9 | 36.7 | 31.2 | 26.7 | 16.1 | 24.2 | 31.2 | 22.7 | 23.6 |
| + RLVR | 41.8 | 41.0 | 50.0 | 30.2 | 42.7 | 40.3 | 49.2 | 26.0 | ||
| + SPELL | 40.0 | 39.2 | 50.9 | 32.3 | 39.9 | 40.1 | 50.8 | 28.2 | ||
| Qwen2.5-14B | 38.0 | 37.2 | 41.9 | 32.1 | 37.3 | 36.2 | 37.5 | 43.3 | 27.5 | 36.1 |
| + RLVR | 52.2 | 51.0 | 63.3 | 32.9 | 53.2 | 52.1 | 64.2 | 30.5 | ||
| + SPELL | 57.6 | 52.6 | 63.0 | 33.5 | 56.8 | 53.0 | 63.2 | 31.2 | ||
| Qwen2.5-32B | 46.8 | 42.6 | 49.0 | 33.7 | 43.0 | 40.7 | 42.2 | 50.1 | 28.7 | 40.4 |
| + RLVR | 58.3 | 50.0 | 59.5 | 32.8 | 57.5 | 49.9 | 60.1 | 32.7 | ||
| + SPELL | 61.8 | 50.2 | 62.1 | 34.2 | 60.6 | 52.2 | 62.3 | 34.3 | ||
| Instruct Models | ||||||||||
| Qwen2.5-7B-Instruct | 38.4 | 40.3 | 45.1 | 29.0 | 38.2 | 39.4 | 41.4 | 44.5 | 28.4 | 38.4 |
| + RLVR | 45.0 | 48.7 | 59.6 | 30.1 | 44.1 | 48.6 | 57.4 | 28.2 | ||
| + SPELL | 45.8 | 46.7 | 63.1 | 33.2 | 44.5 | 48.2 | 60.7 | 32.4 | ||
| Qwen2.5-14B-Instruct | 56.3 | 51.6 | 63.0 | 32.2 | 50.8 | 56.7 | 52.4 | 64.2 | 36.6 | 52.5 |
| + RLVR | 56.1 | 59.6 | 71.0 | 36.4 | 56.7 | 59.9 | 73.4 | 38.5 | ||
| + SPELL | 59.6 | 62.1 | 72.8 | 36.8 | 60.1 | 63.9 | 74.8 | 40.1 | ||
| Qwen2.5-32B-Instruct | 60.0 | 49.9 | 61.4 | 36.0 | 51.8 | 63.0 | 49.4 | 61.5 | 36.2 | 52.5 |
| + RLVR | 59.9 | 60.5 | 70.4 | 36.3 | 59.7 | 62.3 | 69.6 | 36.9 | ||
| + SPELL | 62.3 | 61.2 | 74.4 | 40.1 | 63.3 | 62.0 | 74.1 | 40.8 | ||
| Qwen3-30B-A3B-Instruct | 62.3 | 55.3 | 70.5 | 36.9 | 56.3 | 63.0 | 57.8 | 70.3 | 44.1 | 58.8 |
| + RLVR | 62.5 | 59.9 | 71.8 | 39.8 | 64.0 | 62.0 | 72.4 | 47.4 | ||
| + SPELL | 63.0 | 63.1 | 75.1 | 41.5 | 64.9 | 63.7 | 74.8 | 48.7 | ||
| Llama3.1-8B-Instruct | 33.2 | 45.6 | 52.5 | 29.1 | 40.1 | 34.9 | 47.3 | 53.5 | 27.1 | 40.7 |
| + RLVR | 37.9 | 45.0 | 58.8 | 27.5 | 36.9 | 47.6 | 57.2 | 26.1 | ||
| + SPELL | 39.2 | 48.9 | 61.6 | 28.4 | 39.7 | 50.8 | 60.9 | 26.2 | ||
| Reasoning Models | ||||||||||
| R1-Distill-Llama-8B | 42.0 | 50.3 | 66.8 | 27.9 | 46.8 | 41.5 | 52.6 | 69.3 | 26.4 | 47.5 |
| + RLVR | 43.4 | 51.4 | 67.8 | 30.0 | 45.4 | 54.0 | 68.0 | 28.3 | ||
| + SPELL | 48.9 | 53.4 | 68.4 | 30.2 | 49.2 | 54.3 | 70.0 | 29.3 | ||
| R1-Distill-Qwen-14B | 57.7 | 59.2 | 72.4 | 36.2 | 56.4 | 59.5 | 60.6 | 73.3 | 33.3 | 56.7 |
| + RLVR | 59.6 | 61.7 | 74.6 | 37.2 | 61.0 | 63.8 | 76.0 | 35.9 | ||
| + SPELL | 61.6 | 62.3 | 76.2 | 39.0 | 61.1 | 62.8 | 75.7 | 37.9 | ||
| Qwen3-4B-Thinking | 58.6 | 56.7 | 69.9 | 32.9 | 54.5 | 61.4 | 59.2 | 70.9 | 40.7 | 58.1 |
| + RLVR | 60.5 | 56.6 | 71.1 | 33.8 | 63.3 | 58.6 | 71.1 | 43.4 | ||
| + SPELL | 61.9 | 56.6 | 71.6 | 36.8 | 64.8 | 60.6 | 72.4 | 43.0 | ||
| Qwen3-30B-A3B-Thinking | 62.9 | 64.5 | 75.7 | 39.7 | 60.7 | 63.8 | 65.8 | 77.9 | 46.7 | 63.6 |
| + RLVR | 62.7 | 64.7 | 77.0 | 38.5 | 63.9 | 67.1 | 77.2 | 49.6 | ||
| + SPELL | 64.1 | 66.5 | 78.0 | 42.3 | 66.7 | 68.1 | 78.4 | 50.5 | ||
5.2 Main Results
Table 1 summarizes the results of SPELL across 12 open-source LLMs on six long-context QA benchmarks under maximum input lengths of 16K and 100K tokens. These results offer valuable insights into SPELL’s effectiveness and generalization, as elaborated below.
SPELL consistently enhances performance across diverse models.
Our self-play framework exhibits strong universality, and it delivers substantial improvements across different architectures, sizes, and families. This versatility is evident across the following dimensions. (1) Model types and sizes: SPELL cultivates complex reasoning skills from scratch. For unaligned base models, the average improvement at 16K is large and robust, with Qwen2.5-7B, Qwen2.5-14B, and Qwen2.5-32B improving by 13.9, 14.4, and 9.1 points, respectively. Remarkably, these trained models consistently outperform their instruction-tuned counterparts of the same size, which are trained with extensive human-annotated data. This result highlights that SPELL is data-efficient and practically valuable in scenarios where labeled data is scarce. SPELL also benefits instruction-tuned models, e.g., Qwen2.5-7B-Instruct improves by 9.0 points. For highly specialized reasoning models such as R1-Distill-Qwen-14B, the performance still increases by 3.4 points. (2) Architecture: Beyond dense models, the framework is also applicable to Mixture-of-Experts (MoE) models, where it improves Qwen3-30B-A3B-Instruct and Qwen3-30B-A3B-Thinking by 4.4 and 2.0 points, respectively. (3) Model families: Improvements extend across families. For example, Llama-3.1-8B-Instruct and R1-Distill-Llama-8B increase by 4.4 and 3.4 points, respectively. Collectively, these results establish SPELL as a broadly effective paradigm for advancing LLMs in long-context tasks.
SPELL is superior to traditional RL with static data.
We compare SPELL against the RLVR baseline trained on a fixed dataset synthesized by DeepSeek-R1-0528. Although such static data offers high-quality supervision for RL training, it cannot adapt to the policy’s evolving capabilities. In contrast, SPELL constructs a self-play curriculum that tracks the model’s current ability: the questioner focuses on instances near the responder’s competence boundary, maintaining alignment between the training signal and the policy throughout optimization. The advantage becomes increasingly evident as the policy model’s capabilities grow. For Qwen2.5-7B, RLVR achieves performance comparable to SPELL, indicating that a static corpus appears sufficient for weaker policies. However, for Qwen3-30B-A3B-Thinking, SPELL improves average scores by 2.0, whereas RLVR yields no gain. On the more challenging benchmarks for the same model, RLVR decreases accuracy on DocMath (-0.2) and LongBench-V2 (-1.2), whereas SPELL delivers consistent gains of 1.2 and 2.6 points, respectively. These results validate that when models approach or surpass the quality of static training data, a self-play curriculum proves more effective for sustaining performance gains.
SPELL generalizes to longer contexts.
All models are trained with a 16K input limit and evaluated at 100K without additional tuning. The results remain consistent under this out‑of‑distribution input length, demonstrating that the benefits of SPELL extend beyond the training window. For Qwen2.5‑14B, the average improvement is 14.4 at 16K and increases to 15.0 at 100K. This consistency suggests that the framework strengthens document‑grounded reasoning in a way that remains effective as input lengths grow substantially, rather than producing gains limited to a specific context length.
SPELL boosts exploration and raises the performance ceiling.
We assess test-time exploration with the pass@k metric at a 100K input limit. As shown in Figure 3, Qwen3-30B-A3B-Thinking trained with SPELL exhibits a markedly steeper improvement curve as k increases compared to both the base model and the RLVR baseline. Its pass@8 score reaches 74.5, significantly outperforming the RLVR baseline (68.1) and the original base model (66.9). This enhanced exploratory ability further allows the SPELL-trained model to surpass the performance of the leading gemini-2.5-pro (comanici2025gemini) at a pass@4 rate. These results indicate that SPELL effectively broadens the model’s test-time search space and raises its attainable performance ceiling, highlighting a promising path toward elevating the capabilities of even more powerful foundation models.
5.3 Ablation Studies
To validate the key design choices within the SPELL framework, we conduct ablation studies on Qwen2.5-7B-Instruct. We individually remove each core component of the questioner and verifier roles to quantify their individual contributions to the overall performance.
| Method | DocMath | Frames | LB-MQA | LB-V2 | Average |
| SPELL | 45.8 | 46.7 | 63.1 | 33.2 | 47.2 |
| Questioner | |||||
| w/o Format Penalty | |||||
| w/o Grounding Filter | |||||
| w/o Update | |||||
| w/o History Memory | |||||
| Verifier | |||||
| w/o Verifier | |||||
| w/o Update | |||||
| w/o Majority Voting | |||||
| w/o Update Consistency | |||||
Questioner
As shown in Table 2, the removal of the format penalty and the grounding filter degrades the average score by 1.1 and 1.0 points, respectively. The format penalty keeps the question well-formed, and the grounding filter prevents the generation of hallucinated questions. The largest drops come from disabling the update mechanism and the history memory: freezing the questioner lowers the average score by 4.6, and removing history memory lowers it by 2.9. The declines appear across DocMath, Frames, LB-MQA, and LB-V2, which indicates that these components have a broad impact rather than task-specific effects.
We further examine how these components affect generated question difficulty over Qwen2.5-7B-Instruct training steps, as measured by 1-pass@1 with an external responder (Qwen3-30B-A3B-Instruct) and an external verifier (gpt-oss-120b). The full SPELL model (Figure 4, left) shows a clear upward trend in overall question difficulty, which ensures the questioner proposes questions that are challenging enough for the responder’s evolving capabilities. In contrast, freezing the questioner causes difficulty to stagnate (Figure 4, middle), while removing the history memory makes it erratic (Figure 4, right). The evidence supports the conclusion that continual updates and access to recent history are necessary to form a stable and progressively more challenging curriculum for the responder, which is essential for sustained improvement in a self-play system. This dynamic prevents one role from exploiting the static weaknesses of another, as observed in liu2025spiral.
Verifier
Removing the verifier and relying solely on rule-based rewards decreases average score by 3.2 points, with a 6.4-point drop on DocMath. The CEM-based reward function is brittle and can penalize semantically correct but lexically different answers; the verifier provides a complementary signal in such cases. Interestingly, disabling verifier updates or switching to single-pass decisions leads to moderate declines, which indicates that Qwen2.5-7B-Instruct is already competent at the simpler verification task. However, removing the consistency update mechanism still causes a 1.5-point performance drop. This result shows that the verifier’s updates are susceptible to noise from its own erroneous majority votes, which degrades its reliability. On rule-verifiable tasks, the verifier learns to filter this noise by aligning its majority vote with the ground-truth rule-based outcome. This process provides the verifier with reliable learning signals, which in turn enhance its ability to generate stable rewards for rule-unverifiable outputs. This illustrates how verifiable rewards can guide the calibration of non-verifiable rewards, a finding that aligns with the self-judging methodology in Kimi-K2 (team2025kimi).
5.4 Analysis of Questioner Reward Mapping
We compare our Gaussian-mapped reward function for questioner in Eq. (7) with the reward mapping used in AZR (arz) and R-Zero (rzero). Figure 5(a) visualizes these distinct mapping functions. While the AZR reward function also penalizes high-accuracy questions, it is susceptible to noise from spurious correctness, which can destabilize the training process. In contrast, our Gaussian function, which peaks when the average responder accuracy is 0.5, selectively encourages questions at the frontier of the responder’s competence. Additionally, this mechanism mitigates the impact of data noise. Questions with wrong reference answers typically result in success rates near zero or one, corresponding to scenarios of random guessing on unsolvable questions or consistent matching with the incorrect reference, respectively. Both extremes naturally fall into the low-reward tails of the Gaussian function, effectively suppressing incorrect questions during policy optimization. While R-Zero also centers its peak reward at 0.5, our Gaussian mapping provides a more targeted reward by offering stronger incentives for questions of moderate difficulty and imposing a steeper penalty on those that are either too easy or too hard. This creates a focused and smooth reward distribution that guides the questioner away from generating both trivial and overly difficult questions. The training dynamics corroborate these design differences. As shown in Figures 5(b) and (c), our method maintains a more stable training entropy and exhibits more controlled growth in response length under the Gaussian mapping than under AZR or R-Zero. These advantages in training stability lead to superior overall performance. As Figure 5(d) demonstrates, our method not only achieves more consistent performance growth but also reaches the highest final pass rate among all compared approaches. This evidence supports the view that concentrating the questioner’s reward at the responder’s competence frontier stabilizes the optimization process while preserving headroom for their mutual co-evolution.
5.5 Hyperparameters Analysis
| Settings | DocMath | Frames | LB-MQA | LB-V2 | Average |
| SPELL | 45.8 | 46.7 | 63.1 | 33.2 | 47.2 |
| Group Size () | |||||
| 44.3 | 47.1 | 62.5 | 31.8 | 46.4 | |
| 46.0 | 47.4 | 62.7 | 31.9 | 47.0 | |
| Standard Deviation () | |||||
| 45.3 | 47.0 | 62.3 | 32.8 | 46.9 | |
| 45.5 | 46.0 | 59.4 | 31.8 | 45.7 | |
Sensitivity of group size ()
We examine the impact of the rollout group size on model performance using Qwen2.5-7B-Instruct. As shown in Table 3, while yields the best overall results, SPELL remains robust across different group sizes. We select as the default setting to strike a balance between performance gains and computational efficiency during training.
Selection of standard deviation
The choice of in Eq. (7) is derived from the statistical properties of the Gaussian distribution, where approximately 99.7% of data points fall within three standard deviations () of the mean. In SPELL, we aim to concentrate the questioner’s reward at the point of the responder’s maximal learning efficiency, where . Accordingly, the mean is set to . Given this mean, the distance to either boundary of the valid average responder reward range is 0.5. By setting , we ensure the effective range of the questioner reward covers the responder reward space, yielding . To further validate this theoretical choice, we conduct an ablation study on using Qwen2.5-7B-Instruct. As shown in Table 3, narrowing the curve () has a minimal negative impact, as the reward remains well-focused. However, widening the curve () significantly degrades performance, likely because it assigns higher rewards to overly easy or hard questions, providing a less targeted training signal. This confirms that is both theoretically sound and empirically effective.
5.6 Role of External Judges in Verification
We investigate whether replacing the rule-based judge (CEM-based reward function) with a stronger external model (gpt-oss-120b) benefits the self-play process. As shown in Table 4, introducing a stronger external judge does not yield a significant overall improvement. This suggests that Qwen2.5-7B-Instruct is already capable of learning semantic verification through self-play without external supervision. Notably, when an external judge is introduced, the internal verifier becomes less important; removing it results in only a minor 0.5-point drop, compared to the significant 3.2-point drop observed when using the rule-based judge. This highlights the critical role of the internal verifier in complementing the brittle CEM-based reward function when an external judge is not available.
| Method | DocMath | Frames | LB-MQA | LB-V2 | Average |
| Qwen2.5-7B-Instruct | 38.4 | 40.3 | 45.1 | 29.0 | 38.2 |
| + SPELL (Rule-based Judge) | 45.8 | 46.7 | 63.1 | 33.2 | 47.2 |
| + SPELL (Gpt-oss-120b Judge) | 47.1 | 48.0 | 61.6 | 32.1 | 47.2 |
| + SPELL (Rule-based Judge) w/o Verifier | 39.4 | 46.6 | 60.4 | 29.4 | 44.0 |
| + SPELL (Gpt-oss-120b Judge) w/o Verifier | 47.0 | 47.2 | 61.1 | 31.3 | 46.7 |
6 Conclusion
This work introduces SPELL, a multi-role self-play reinforcement learning framework for evolving the long-context reasoning capabilities of LLMs without human supervision. A single policy model alternates among the roles of questioner, responder, and verifier to generate questions, solve them, and assess the solutions, which reduces reliance on costly and unreliable human annotation while enabling stable self-evolution. Extensive experiments across 12 models of diverse architectures and sizes show that SPELL delivers consistent and substantial improvements in long-context reasoning.
This study concludes with three notable findings. First, signals from verifiable tasks can calibrate and strengthen the verifier’s assessment on non-verifiable outputs, thereby ensuring a reliable self-rewarding mechanism. Second, within a multi-role self-play framework, sustaining a dynamic equilibrium among the capabilities of different roles is critical for the stable evolution of the shared policy. Finally, our results demonstrate that for models approaching or surpassing human performance, where external supervision emerges as a fundamental bottleneck, autonomous self-evolution transitions from a promising alternative to an indispensable strategy for sustained advancement.
Ethics Statement
This research focuses on the development of long-context LLMs through self-play that requires no human supervision. While we believe our methodology does not inherently raise significant ethical issues, we acknowledge the potential for misuse of this technology. We also recognize that an unsupervised learning approach may perpetuate or amplify societal biases in the model. Our research is conducted using only publicl y available datasets, in compliance with their licenses, and involves no personally identifiable information. We have adhered to all relevant ethical and legal standards and declare no conflicts of interest that could have influenced the outcomes of this study.
Reproducibility Statement
To ensure the reproducibility of our work, we provide full experimental details in Section 5.1 and Appendix C. These include our methods for dataset construction, training configurations, and the evaluation setup. The implemented code, the data used, and a comprehensive guide to reproduce our method are available in the supplementary materials.
Appendix A Limitations and Future Work
First, while our study provides strong empirical evidence, a theoretical framework explaining the co-evolutionary dynamics of the three roles within a single model has yet to be explored. Second, due to framework and efficiency constraints, our experiments are limited to a 16K input context. Although the acquired skills generalize well to longer contexts, it is a critical next step to develop more efficient frameworks for self-play RL on ultra-long contexts, such as 128K tokens and beyond. Finally, our self-play framework still relies on a degree of human intervention, such as pre-processing the document corpus and crafting prompt templates for each role. Future work could explore pathways toward greater autonomy, such as a system where an LLM interacts with a real-world environment to generate and evolve its own tasks, templates, and reward functions.
Appendix B SPELL Algorithm
In this section, we outline the step-by-step algorithm for SPELL in Algorithm 1.
Appendix C Implementation Details
C.1 Training Data Construction
Our training data supports three distinct tasks: document general QA, financial math QA, and multiple-choice QA. We construct the dataset from two complementary sources. The first is the DocMath dataset (zhao2024docmath), which provides specialized data comprising long, complex financial reports that necessitate numerical reasoning. We used only the raw documents, discarding the original unlabeled questions. From this dataset, we select 2,150 instances with a total token length below 16K, which are designated for the financial math QA task.
The second component is a general-domain corpus designed to enhance the model’s fundamental document understanding. This dataset originates from the 1.16 billion English documents in the Ultra-Fineweb corpus (ultrafineweb) and is curated through the following multi-stage pipeline. First, an initial filtering stage retains high-quality texts of appropriate length by selecting documents with a perfect quality score of 1.0 and character lengths between 100 and 32,768. This step downsamples the corpus to a high-quality subset of 4 million documents. Second, a cleaning procedure ensures data diversity and prevents test set contamination. We employ the MinHashLSH algorithm (minhash) with a threshold of 0.8 and 128 permutation functions for deduplication and to decontaminate the data against all documents in our evaluation benchmarks. This cleaning phase refines the corpus to 1.1 million unique and decontaminated texts. Third, a hierarchical clustering approach structures the dataset into different topics. We generate 4096-dimensional embeddings for each text using the Qwen3-Embedding-8B model (qwen3_embedding), which are then grouped into 50,000 distinct clusters via a hierarchical k-means algorithm (xu2015hierarchical). Each cluster contains an average of 20 semantically related documents, and the resulting domain distribution, visualized in Figure 6 (left), confirms the dataset’s broad topical diversity. This curated corpus from Ultra-Fineweb is used to generate training instances for the document general QA and multiple-choice tasks. To ensure a comparable scale across tasks, the number of clusters selected for these two tasks matches the data size of the DocMath portion. Ultimately, our combined training data consists of 6,450 unique document sets, with 2,150 designated for each of the three tasks.
C.2 Evaluation Details
Evaluation Benchmarks
We evaluate our models using a suite of well-established benchmarks designed to assess long-context comprehension and reasoning. These benchmarks fall into two primary categories: multiple-choice and multi-hop question answering (QA). For the multiple-choice task, we use LongBench-V2 (longbench_v2), a benchmark of 503 questions that assesses deep comprehension across six areas: single-document QA, multi-document QA, long in-context learning, long-dialogue history understanding, code repository understanding, and long structured data understanding. For multi-hop QA, our evaluation incorporates several benchmarks: Frames (krishna2024fact), containing 824 questions on diverse Wikipedia topics such as history, sports, science, animals, and health; three subsets from LongBench (longbench_v1), each with 200 questions, namely HotpotQA (yang2018hotpotqa) (2-hop), 2WikiMultihopQA (ho2020constructing) (requiring up to five hops), and MuSiQue (trivedi2022musique) (requiring up to four hops); and DocMath (zhao2024docmath), which focuses on numerical reasoning within financial reports. For DocMath, we use the testmini subset of 800 queries, which is orthogonal to our training data. As shown in Figure 6 (right), the test instances across these benchmarks cover a wide range of context lengths.
Evaluation Configurations
We evaluate all models at two maximum input lengths: 16K tokens, which aligns with our training configuration, and 100K tokens to test for generalization to longer contexts. The maximum generation length is 4K tokens for non-reasoning models and is extended to 20K tokens for reasoning models. For prompts exceeding the maximum context window, we employ the middle truncation strategy from longbench_v1 to preserve the front and tail portions of the context. All experiments are conducted using a sampling temperature of 0.7 and a top- value of 0.95. For each query, we generate candidate responses, reporting the average score (pass@1) for our main experiments and the pass@k metric for the test-time scaling analysis. The pass@k metric is an unbiased estimator for the probability that at least one of candidate solutions is correct, given candidates per problem, of which are correct. It is calculated as:
Our scoring is tailored to each benchmark’s format. For multiple-choice tasks, we report standard accuracy. For multi-hop QA tasks, simple string matching is often insufficient to assess the semantic correctness of free-form text answers. Thus, we supplement the cover exact match (CEM) score with LLM-as-a-judge (zheng2023judging), which uses gpt-oss-120b (gpt_oss_system_card) to evaluate semantic equivalence between a model’s prediction and the ground-truth answer. The prompt for this evaluation is detailed in Table 5. The final score for these tasks is the maximum of the two, providing a more comprehensive and robust assessment of model performance.
C.3 RL Algorithm Details
To enhance stability and practical performance, we integrate two key techniques inspired by Decoupled Clip and Dynamic Sampling Policy Optimization (DAPO) (yu2025dapo) for the baselines and our method. First, we adopt a token-level policy gradient loss, which normalizes each token’s contribution by the total number of tokens in the group. This approach ensures every token in the same group contributes equally to the final objective, which prevents the learning signal from valuable tokens in high-quality, long responses from being diluted while ensuring that undesirable patterns in low-quality, lengthy outputs are effectively penalized. Second, we employ a dynamic sampling strategy: any group of generations where the rewards exhibit zero variance is discarded from the training batch. This ensures the advantage in Eq. (3) is always well-defined.
Consistent with recent findings suggesting that removing KL regularization can improve exploration and accelerate convergence (OpenReasonerZero2025; yu2025dapo; qwenlongl1), we set . Besides, we operate in a strictly on-policy setting, performing only a single gradient update per batch of samples. This design choice implies that the policy being updated, , remains identical to the policy that generated the data, . Since the importance sampling ratio is strictly equal to 1, the clipping function becomes inactive, and we can remove it from the objective. Note that the advantage is independent of , the training objective in Eq. (2) simplifies to:
C.4 Details of Baselines
We compare SPELL with two categories of baselines. The first category comprises the base models on which SPELL is built, organized into three groups: (1) base models, including Qwen-2.5-7B, Qwen-2.5-14B, and Qwen-2.5-32B; (2) instruction-tuned models, including Qwen-2.5-7B-Instruct, Qwen-2.5-14B-Instruct, Qwen-2.5-32B-Instruct, Qwen3-30B-A3B-Instruct, and Llama-3.1-8B-Instruct; and (3) reasoning models, including R1-Distill-Llama-8B, R1-Distill-Qwen-14B, Qwen3-4B-Thinking, and Qwen3-30B-A3B-Thinking (qwen3_technical_report; yang2024qwen2; guo2025deepseek; dubey2024llama). The second baseline is traditional RLVR. For data construction, we use DeepSeek-R1-0528 (guo2025deepseek) as both the questioner and responder to synthesize a dataset from the same document clusters used for SPELL. The synthesized dataset is then filtered by a verifier model (gpt-oss-120b), and retains only the instances where the answers from the questioner and responder are identical. This process yields a dataset of approximately 3,000 verifiable samples that covers multiple-choice, document general QA, and financial math QA tasks. For RLVR training, we employ the cover exact match (CEM) as the rule-based reward function. We generate eight trajectories per question and maintain all other hyperparameters identical to those of SPELL to ensure a fair comparison.
C.5 Details of Open-Source Models and the Dataset
In Table 6, we provide the Huggingface repository names of all policy models, the embedding model, the judge model, and datasets used in our experiments.
| Model/Dataset | Huggingface ID |
| Qwen2.5-7B | Qwen/Qwen2.5-7B |
| Qwen2.5-14B | Qwen/Qwen2.5-14B |
| Qwen2.5-32B | Qwen/Qwen2.5-32B |
| Qwen2.5-7B-Instruct | Qwen/Qwen2.5-7B-Instruct |
| Qwen2.5-14B-Instruct | Qwen/Qwen2.5-14B-Instruct |
| Qwen2.5-32B-Instruct | Qwen/Qwen2.5-32B-Instruct |
| Qwen3-30B-A3B-Instruct | Qwen/Qwen3-30B-A3B-Instruct-2507 |
| Llama-3.1-8B-Instruct | meta-llama/Meta-Llama-3.1-8B-Instruct |
| R1-Distill-Llama-8B | deepseek-ai/DeepSeek-R1-Distill-Llama-8B |
| R1-Distill-Qwen-14B | deepseek-ai/DeepSeek-R1-Distill-Qwen-14B |
| Qwen3-4B-Thinking | Qwen/Qwen3-4B-Thinking-2507 |
| Qwen3-30B-A3B-Thinking | Qwen/Qwen3-30B-A3B-Thinking-2507 |
| Qwen3-Embedding-8B | Qwen/Qwen3-Embedding-8B |
| gpt-oss-120b | openai/gpt-oss-120b |
| DocMath | yale-nlp/DocMath-Eval |
| Ultra-Fineweb | openbmb/Ultra-FineWeb |
Appendix D Additional Analysis
This section provides further insights into the properties of SPELL. We begin by analyzing the training cost in Section D.1 and exploring generalization to short-context tasks in Section D.2. We then extend our evaluation to additional long-context benchmarks in Section D.3 and compare our approach with state-of-the-art long-context alignment baselines in Section D.4. The section concludes with an analysis of the training dynamics of role evolution in Section D.5 and a discussion on reward hacking risks and mitigations in Section D.6.
D.1 Training Cost Analysis
We analyze the computational cost of SPELL using the Qwen2.5-7B-Instruct model on a single node with 8 80GB NVIDIA A100 GPUs. Figure 7 illustrates the time breakdown for the two primary stages of our framework: role-specific rollout (questioning, responding, and verifying) and unified updating. The total time per training step averages approximately seven minutes, and the total training cost is about eight hours. Although the number of verifier rollouts is times greater than that of the responder, as outlined in Section 4.3, its computational cost is lower. This efficiency stems from two factors: the verifier processes shorter inputs, as it does not require the long documents provided to the responder, and its task of generating brief verifications is less demanding than the responder’s task of reasoning over long contexts. During the later stages of training, the time for verifying constitutes only about half of the time required for responding. Consequently, generating verification judgments for each response to create a reliable reward signal does not introduce a significant computational bottleneck. As shown in Section 5.3 and Table 2, the verifier provides a significant 3.2-point performance gain (from 44.0 to 47.2). The significant performance gain by introducing the verifier far outweighs the minor increase in training cost.
D.2 Short-Context Reasoning Results
We further investigate the generalization of our proposed SPELL method from long-context reasoning to short-context reasoning tasks. Our evaluation suite includes five mathematical benchmarks and the MMLU Pro general knowledge benchmark as follows:
AIME 24/25 (AIME) is the American Invitational Mathematics Examination, a prestigious high-school mathematics competition administered by the Mathematical Association of America (MAA). The AIME consists of two exams (I and II) annually, each containing 15 problems, for a total of 30 problems per year. The answer to each question is an integer from 0 to 999. These problems demand deep mathematical knowledge and creative problem-solving strategies, making them a challenging benchmark for advanced mathematical reasoning.
AMC 23444https://huggingface.co/datasets/AI-MO/aimo-validation-amc (yang2024qwen25math) refers to the 2023 American Mathematics Competition. This competition features 25 questions designed to test advanced high school mathematics, covering topics such as trigonometry, advanced algebra, and advanced geometry.
MATH (hendrycks2021math) is a dataset of math problems ranging in difficulty from middle school to high school competition level. It is designed to test a wide range of mathematical skills, including algebra, geometry, number theory, and counting & probability. For our evaluation, we utilize a subset of 500 problems, referred to as MATH-500.
GSM8K (cobbe2021gsm8k) is a dataset of 1,319 grade school math word problems. These problems are designed to be solvable by a capable middle school student and require two to eight steps of reasoning using basic arithmetic operations.
MMLU-Pro (wang2024mmlupro) is an enhanced version of the MMLU (hendrycks2021mmlu) dataset, designed to address issues such as noisy data and reduced difficulty resulting from advances in model capabilities and increased data contamination. MMLU-Pro raises the difficulty by expanding the multiple-choice options from four to as many as ten and incorporating expert-reviewed annotations for improved quality and reduced noise.
| Model | AIME24 | AIME25 | AMC23 | MATH500 | GSM8K | MMLU Pro | Average |
| Qwen2.5-7B | 5.42 | 3.33 | 33.44 | 53.60 | 76.57 | 40.24 | 35.43 |
| + SPELL | |||||||
| Qwen2.5-14B | 6.67 | 5.83 | 37.81 | 61.68 | 84.31 | 46.67 | 40.50 |
| + SPELL | |||||||
| Qwen2.5-32B | 9.17 | 5.83 | 45.31 | 66.25 | 87.34 | 48.89 | 43.80 |
| + SPELL |
Consistent with our main experiments, all models are evaluated with a maximum output of 4K tokens, a sampling temperature of 0.7, and a top- value of 0.95. We report the accuracy averaged over 8 independent runs for each task. As shown in Table 7, SPELL, trained solely on the long-context data, improves performance on short-context reasoning benchmarks across all base models. The average scores of Qwen2.5-7B, Qwen2.5-14B, and Qwen2.5-32B increase by 6.93, 8.74, and 7.24 points, respectively. The consistent gains indicate that reasoning competencies acquired through long-context self-play transfer effectively to short-context settings.
D.3 Additional Long-Context Benchmarks
We further evaluate SPELL on two challenging long-context benchmarks: MRCR (MRCR)555https://huggingface.co/datasets/openai/mrcr, a multi-needle “Needle in a Haystack” benchmark, and three subsets of HELMET (HELMET), which covers Retrieval-Augmented Generation (RAG), In-Context Learning (ICL), and Summarization (Summ). We evaluate the Qwen2.5 base models against the RLVR baseline and SPELL with a maximum input length of 16K and maximum output length of 4K. The results in Table 8 show that SPELL consistently and significantly outperforms both the base models and the RLVR baseline across these diverse long-context tasks, demonstrating strong generalization capabilities beyond the standard QA tasks.
| Model | MRCR-2needle | MRCR-4needle | MRCR-8needle | Helmet-RAG | Helmet-ICL | Helmet-Summ | Average |
| Qwen2.5-7B | 6.9 | 2.0 | 2.2 | 50.0 | 3.5 | 4.1 | 11.5 |
| + RLVR | 22.0 | 12.5 | 10.5 | 49.3 | 2.6 | 14.3 | 18.5 |
| + SPELL | 34.5 | 16.5 | 16.0 | 54.2 | 10.4 | 13.7 | 24.2 |
| Qwen2.5-14B | 23.1 | 10.5 | 10.0 | 42.4 | 1.5 | 3.7 | 15.2 |
| + RLVR | 20.9 | 9.5 | 9.1 | 46.7 | 42.0 | 24.7 | 25.5 |
| + SPELL | 35.0 | 22.1 | 17.8 | 52.3 | 39.2 | 23.0 | 31.6 |
| Qwen2.5-32B | 27.0 | 11.5 | 12.0 | 59.0 | 42.8 | 16.2 | 28.1 |
| + RLVR | 36.7 | 14.6 | 13.0 | 52.7 | 16.7 | 21.0 | 25.8 |
| + SPELL | 38.0 | 18.5 | 14.7 | 61.4 | 56.4 | 21.2 | 35.0 |
D.4 Comparison with Long-Context Alignment Baselines
We compare SPELL against three recent long-context alignment baselines—SoLoPO (solopo), LongPO (longpo), and QwenLong-L1 (qwenlongl1)—using Qwen2.5-7B-Instruct as the base model. To ensure a fair comparison, we reimplement these methods using the same document corpus employed in SPELL. For SoLoPO and LongPO, the core step is to construct preference pairs from short contexts containing key information and long contexts containing distractors given the same question. Specifically, for each data instance comprising documents, we first randomly sample documents as the short text and employ DeepSeek-R1-0528 as the questioner to generate QA pairs. Then, we use DeepSeek-R1-0528 as the responder to answer the proposed questions using the full set of documents. We retain only those QA pairs where the answers from the questioner and responder are consistent. Next, we take the 5 documents from the questioning stage as short texts, corresponding to in SoLoPO and in LongPO, and take all documents as long texts, corresponding to in SoLoPO and in LongPO. Finally, we apply their respective preference pair construction strategies and training configurations on Qwen2.5-7B-Instruct to reproduce these methods. For QwenLong-L1, we use
| Model | DocMath | Frames | LB-MQA | LB-V2 | Average |
| Qwen2.5-7B-Instruct | 38.4 | 40.3 | 45.1 | 29.0 | 38.2 |
| + RLVR | 45.0 | 48.7 | 59.6 | 30.1 | 45.9 |
| + LongPO (Reimpl.) | 41.4 | 44.2 | 53.7 | 32.0 | 42.8 |
| + LongPO (Official) | 42.3 | 41.4 | 45.7 | 30.9 | 40.1 |
| + SoLoPO (Reimpl.) | 45.3 | 43.9 | 56.0 | 31.6 | 44.2 |
| + QwenLong-L1 (Reimpl.) | 45.6 | 46.7 | 60.0 | 32.0 | 46.1 |
| + SPELL (Ours) | 45.8 | 46.7 | 63.1 | 33.2 | 47.2 |
the same synthesized data as our RLVR baseline and follow their official GRPO training setup. We evaluate both our reimplemented models and the official LongPO checkpoint using a maximum input length of 16K and a maximum output length of 4K. The results in Table 9 demonstrate that SPELL consistently outperforms these long-context alignment baselines.
D.5 Evolutionary Dynamics of Questioner and Verifier
To understand the self-evolutionary process and identify potential failure modes, we analyze the behavior of the questioner and verifier roles across different training steps.
Questioner dynamics
We track the distribution of valid questions generated by the questioner throughout the training process. As shown in Figure 8(a), the task distribution is notably imbalanced during the first 10 steps, with Financial Math QA accounting for over 70% of solvable tasks. This imbalance likely occurs because the model transfers its strong mathematical reasoning capabilities to the Financial Math QA task, which necessitates substantial numerical calculation. As training progresses and the policy evolves, the distribution becomes increasingly balanced. This indicates that SPELL effectively drives the questioner to explore a diverse range of task types.
Verifier calibration
We analyze the disagreement rate between the verifier’s majority vote and the rule-based judge (CEM). Figure 8(b) plots this metric for General QA and Financial Math QA, both of which contain questions that are partially non-verifiable by strict rules. Initially, the verifier struggles with General QA, which often involves open-ended semantic equivalence that rule-based checks fail to capture. In contrast, Financial Math QA exhibits a lower initial disagreement rate, attributed to the model’s relatively strong numeric reasoning ability. The disagreement rate consistently decreases for both tasks, indicating that the verifier progressively aligns with the rule-based judge. This trend further suggests that the verifier’s updates guide the questioner toward generating questions that are more verifiable by the rule-based judge.
D.6 Analysis of Reward Hacking
Reward hacking is a significant concern in self-play systems. Throughout our exploratory experiments, we identified several potential failure modes and implemented specific mitigations.
Questioner stagnation
Without the automated curriculum and history memory, the questioner tends to repeatedly propose similar, trivial questions about the same document to maximize the responder’s success rate. The history memory module conditions the questioner on recently solvable QA pairs and newly introduced documents. The prompt in Appendix E explicitly instructs the questioner to produce novel and more difficult questions.
Responder mode collapse
When the responder receives rewards for outputs that are merely substrings of the ground-truth answer, it can hack the metric by generating short phrases like “The answer is”. We address this by enforcing a stricter cover exact match (CEM) criteria, which requires the responder to include the complete ground-truth answer to receive a positive reward.
Verifier self-delusion
Updating the verifier solely based on its own majority vote as a pseudo-label can lead to verifier hacking. Without external supervision, verification errors accumulate, eventually causing the verifier to label all responders’ outputs as correct. To mitigate this, we introduce the consistency check mechanism, which aligns the verifier’s judgments with rule-based rewards on verifiable questions, thereby preventing this self-delusion.
These observations confirm that the architectural components of SPELL—specifically the history memory, prompt templates, consistency checks, and CEM-based reward function—are essential to mitigate reward hacking and ensure stable self-evolving.
Appendix E Prompt Template
In this section, we detail the prompt templates for the questioner, responder, and verifier across all tasks. For the questioner, we apply different prompting strategies for document clusters with and without history memory; these prompts are modified from the guidelines for human annotators in LongBench-V2 (longbench_v2). The responder prompt for financial math QA is modified from DocMath (zhao2024docmath), and the prompts for document general QA and multiple-choice are modified from LongBench-V2. The verifier prompt for document QA tasks is modified from Frames (krishna2024fact) and requires the model to evaluate the semantic equivalence of a generated answer against a ground-truth reference.